Field Notes
<--2024-04-25 10:15:00 | 2024-05-23 12:25:00-->Other sites visited today: Gilbert Tract | East End
Phenocam link: Searching...
Mayberry: 2024-05-07 11:05:00 (DOY 128)
Author: Daphne Szutu
Others: Kyle
Summary: Greening up, goats have cleared area by tower, sampled 5 sipper profiles, replaced clogged sippers
2024-05-07 Mayberry Kyle and I arrived at 11:05 PST (12:05 PDT). It was sunny, mild, and breezy. We had a small rainstorm a few days ago, and that's probably the last of the rain this season. There were goats on the levee road by Mayberry Slough. We opened and drove through two panel gates. We saw a livestock guard dog but he wasn’t very aggressive, just came to look at us. There was also a net electric fence up across the interior levee road, but it looked in bad shape and parts of it was ripped. I just drove over a section where a single string ran across the road. The goats have been by our tower and solar panels. Almost everything has been cleared between our usual parking spot and the tower, except for the very top flowers of the hemlock. Mayberry is starting to look more than half green. I mostly see dark green tules for now. The staff gauge read 31cm. Kyle got in the water and started sampling her open water sippers, including the two profiles she just installed last week. All of the sippers that had been developed produced water, yay. Some of the deeper depths (30-50cm) had many bubbles coming up along the tubing. Kyle thinks it’s mostly methane bubbles being generated in the sediment. To sample, Kyle uses a 60ml syringe with three valves, two one-way valves and one three-way valve. She pulls up about 5-15ml of water from the sipper to flush it, then discards that water. She then pulls up another ~30ml to measure using the handheld conductivity sensor. The sensor is not rinsed between samples. I downloaded met, cam, and USB GHG data and cleaned flux and rad sensors. I topped off the wash reservoir. The picam got a little tilted, so I straightened it. I changed the 7700 heater from winter to summer hours. Previously, the heater turned on from 18:00 to 06:00, now the heater will turn on from 20:00 to 06:00. The 7500 read: 433ppm CO2, 382mol/m3 H2O, 20.4C, 101.1kPa, 99SS – 100 after cleaning The 7700 read: 1.97ppm CH4, 54RSSI – 73 after cleaning I then worked on sampling and measuring the veg-blue and the veg-black profiles. The veg-black profile had two sippers that have been clogged for a while; Kyle replaced those with freshly installed sippers and developed them. The veg-red profile is mostly clogged, so Kyle pulled out all of the sippers and did a fresh installation of new sippers. She develops them by drawing surface water through them so they were full of water before installing. Then after pushing them into the ground using the installation tool, she uses the syringe to push about 60ml of surface water through the whole sipper. At the end, Kyle noticed there were two red-red sippers—-one must be the old one she forgot to remove. She will sample both of them next time and remove the red-red sipper that is clogged (which should be the old sipper). For faster sampling, I tested out the small 12V peristaltic pump that we used to test precip buckets. It worked well to pull up water from the sippers, although since most of the sippers are pretty easy to sample now, I’m not sure the pump will speed things up. Open-red-yellow, 7.9mS, 20.5C Open-red-red, 7.09mS, 18.2C, easy to pull with relatively clear sample and not many bubbles Open-red-white, 7.9mS, 18.8C Open-red-black, 7.81mS, 18.7C Open-red-green, 7.18mS, 19.1C, easy to pull with relatively clear sample and not many bubbles Surface, 1.74mS, 18.2C Open-blue-yellow, 5.44mS, 19.2C Open-blue-red, 6.51mS, 19.0C Open-blue-white, 6.5mS, 18.1C Open-blue-black, 6.15mS, 19.0C Open-blue-green, 5.05mS, 18.1C, easy to pull with relatively clear sample and not many bubbles Open-black-yellow, 7.75mS, 16.7C Open-black-red, 6.72mS, 17.1C Open-black-white, 4.92mS, 16.4C Open-black-black, 4.9mS, 17.6C Open-black-green, 7.23mS, 17.3C, many bubbles Veg-blue-yellow, 11.12mS, 20C Veg-blue-red, 9.84mS, 19.6C Veg-blue-white, 8.9mS, 18.8C Veg-blue-black, 8.8mS, 19.6C, many bubbles Veg-blue-green, 7.44mS, 20.4C Veg-black-yellow, 5.29mS, 19.7C Veg-black-red,,,clogged Veg-black-white,,,clogged Veg-black-black, 9.68mS, 22.0C Veg-black-green, 5.43mS, 19.0C Surface, 1.9mS, 19.6C, after tromping around in the wetland took a conductivity reading where I had been standing We left at 13:00 PST (14:00 PDT). On the way out, there was a new net electric fence crossing over the ramp down from the levee, just before Antioch bridge. The shepherd was right there and he let us through. He was wondering when we would return, and Kyle conversed with him in Spanish to say we would be back in two weeks. |
1 photos found
20240507MB_KyleSippers.jpg ( 2024-05-07 12:23:32 ) Full size: 1440x1920
Kyle sampling porewater sippers in the channel
11 sets found
7 MB_flux graphs found
No data found for MB_met.
28 MB_tule graphs found
Can't check MBWPEESWGT_mixed
1 MB_cam graphs found
Can't check MBEEWPSWEPGT_cam
Can't check SIMBTWEE_met
1 MB_minidot graphs found
Can't check MBSWEEWPBCBA
1 MB_processed graphs found
Can't check MB_fd
1 MB_fd graphs found